Lee has an older sister and a younger sister. Her older sister is one year more than twice Lee's age. Lee's younger sister is three years younger than she is. The sum of their three ages is 26. Find Lee's age?
Answer by rapture(86)

Your question as typed yields a fractional answer, which is not possible in terms of age. I think you meant to say that Lee's younger sister is three years younger than him not the older sister.

Let x = Lee

Let 2x + 1 = older sister

Let x - 3 = younger sister

The sum of their ages is

x + (2x + 1) + (x - 3) = 26

Solve for x to find Lee's age.